A 0.70 kg ball moving horizontally at 6.0 m/s strikes a vertical wall and rebounds with speed 3.5 m/s.what is the magnitude of the change in its linear momentum.
1
Expert's answer
2013-01-29T09:09:04-0500
Linear momentum is p=m*v, where m is mass and v -velocity Magnitude of change is p1-p2=m1*v1-m2*v2 But mass isn't change, so

p1-p2=m*v1-m*v2=m*(v1-v2)=0,7*(6-3,5)=0,7*2,5=1,75 m*kg/s